OpenStudy (anonymous):

I need some help with this problem and It needs to be solved in Dimensional Analysis form. Fingernails grow 1nm/sec. How many cm do fingernails grow in one year?

OpenStudy (anonymous):

I would start with a year and work your way to seconds and it should leave you with your answer in nm

OpenStudy (anonymous):

\[1 year (\frac{365 days}{1 year})(\frac{24 hours}{1 day})(\frac{60 \min}{1 hour})(\frac{60 \sec}{1 \min})(\frac{1 nm}{1 \sec})(\frac{1x10^-9 m}{1 nm})(\frac{1 cm}{1x10^-2 m})=\]

OpenStudy (anonymous):

Thank you!! This helped alot! I just did not know how to start it.

OpenStudy (anonymous):

In problems like these there should always be a number or quantity that isn't a conversion factor. In this case it was the year, so 9 times out of 10 you are going to start with the one you know the least about or isn't a conversion factor.